# Set and functional prediction: randomness, exchangeability, and conformal

### v1

@@ 3 Main results
Let $B$ be a Markov kernel with source $\mathbf{Z}$ and target $\mathbf{Y}$, which we will write in the form $B:\mathbf{Z}\hookrightarrow\mathbf{Y}$ (as in [10, Sect. A.4]). We will write $B(A\mid z)$ for its value on $z\in\mathbf{Z}$ and $A\subseteq\mathbf{Y}$ (where $A$ is measurable), and we will write $\int f(y)B(\mathrm{d}y\mid z)$ for the integral of a function $f$ on $\mathbf{Y}$ w.r. to the measure $A\mapsto B(A\mid z)$. We will show that the efficiency of various predictors (such as the conformal predictor) derived from a randomness predictor $E$ is not much worse than the efficiency of the original randomness predictor $E$ on average, and $B$ will define the meaning of “on average”.
3.1 Kolmogorov’s step
The following statement shows that the efficiency does not suffer much on average when we move from randomness e-prediction to exchangeability e-prediction. It is a counterpart of Corollary 5 in [8] (with a similar proof).
@@ 3.1 Kolmogorov’s step
(with $/0$ 0 interpreted as 0 and $z_{n+1}$ represented as $(x_{n+1},y_{n+1})$) is a randomness e-variable.
We can interpret (9) as a statement that $E^{\mathrm{X}}$ is almost as efficient as $E$: the mean ratio of the degree to which $E$ rejects a false label $y$ to the degree to which $E^{\mathrm{X}}$ rejects $y$ is bounded by $\mathrm{e}$ under any probability measure that may depend on the test example. This will be further discussed after we state our main result, Theorem 8.
Proof of Theorem 3.
@@
Combining all these statements, we get
$G(z_{1},\dots,z_{n+1})$ $=G_{2}(z_{1},\dots,z_{n+1})G_{3}(z_{1},\dots,z_{n+1})$
$\geq\frac{1}{\mathrm{e}}G_{1}(z_{1},\dots,z_{n+1})G_{3}(z_{1},\dots,z_{n+1})$
$=\int E^{\mathrm{i}}(z_{1},\dots,z_{n},x_{n+1},y)B(\mathrm{d}y\mid z_{n+1}).$
By the definition (6), this is equivalent to (9). ∎
Notice that Theorem 3 does not assume the homoscedasticity of the labels. The simplest informative examples, however, are indeed homoscedastic: in them, for each $z=(x,y)\in\mathbf{Z}$, $B(\cdot\mid z)$ is the distribution of $y+\xi$ for a given random variable $\xi$. In general, however, the distribution of $\xi$ may depend on the object $x$.
The following result is a simple inverse to Theorem 3.
Theorem 4.
The constant $\mathrm{e}^{-1}$ in Theorem 3 cannot be replaced by a larger one.
Proof.
In this proof we follow the example in [8, Sect. B.1] (the example in [8] is informal, and here we formalize it). Without loss of generality we assume $\left|\mathbf{X}\right|=1$ (so that the objects become uninformative and we can omit them from our notation) and $\mathbf{Y}=\{0,1\}$ (with the discrete $\sigma$-algebra). Define a randomness e-variable $E$ by
@@
where $k$ is the number of 1s in $y_{1},\dots,y_{n+1}$. This is indeed a randomness e-variable, since the maximum probability of $k=1$ in the Bernoulli model, $(n+1)p(1-p)^{n}\to\max$, is attained at $p=\frac{1}{n+1}$. The corresponding exchangeability e-variable is
@@
Let $B$ just flip the label: $B(\{1-y\}\mid y)=1$. Suppose Theorem 3 holds with the $\mathrm{e}^{-1}$ in (9) replaced by $c>\mathrm{e}^{-1}$. Then the randomness e-variable $G$ satisfies
@@
which is impossible for a large enough $n$ (since the probability measure concentrated on $(0,\dots,0)$ is of the form $Q^{n+1}$). ∎

### v2

@@ 3 Main results
Let $B$ be a Markov kernel with source $\mathbf{Z}$ and target $\mathbf{Y}$, which we will write in the form $B:\mathbf{Z}\hookrightarrow\mathbf{Y}$ (as in [17, Sect. A.4]). We will write $B(A\mid z)$ for its value on $z\in\mathbf{Z}$ and $A\subseteq\mathbf{Y}$ (where $A$ is measurable) and write $\int f(y)B(\mathrm{d}y\mid z)$ for the integral of a function $f$ on $\mathbf{Y}$ w.r. to the measure $A\mapsto B(A\mid z)$. We will show that the efficiency of the conformal predictor derived from a train-invariant randomness predictor $E$ is not much worse than the efficiency of the original randomness predictor $E$ on average, and $B$ will define the meaning of “on average”.
3.1 From train-invariant randomness e-prediction to conformal e-prediction
The following statement shows that the efficiency does not suffer much on average when we move from randomness e-prediction to exchangeability e-prediction. It is more general than what we need at the moment, since it also covers randomness e-predictors that are not train-invariant.
@@ 3.1 From train-invariant randomness e-prediction to conformal e-prediction
(with $/0$ 0 set to 0 and $z_{n+1}$ represented as $(x_{n+1},y_{n+1})$) is a randomness e-variable.
We can interpret (7) as a statement that $E^{\mathrm{X}}$ is almost as efficient as $E$ unless the true data sequence $z_{1},\dots,z_{n+1}$ does not look IID. The “unless” clause makes sense in view of our assumption of randomness. According to (7) and assuming that $B(\cdot\mid z_{n+1})$ is concentrated on $\mathbf{Y}\setminus\{y_{n+1}\}$, the mean ratio of the degree to which $E$ rejects a false label $y$ to the degree to which $E^{\mathrm{X}}$ rejects $y$ is not large under any probability measure that may depend on the test example unless we can reject the assumption of randomness for the true data sequence. The gap $\mathrm{e}^{-1}\approx 0.37$ between the mean ratio and the e-value at which we reject the assumption of randomness is optimal (Theorem 6 below).
From now on, until the end of the main paper, let us assume that $E$ is train-invariant in Theorem 5, i.e., $E$ is a train-invariant randomness e-predictor (the only exception is the line following the statement of Corollary 7). Then $E^{\mathrm{X}}$ in (7) is a conformal e-predictor.
A full proof of Theorem 5 will be given in Sect. A.1, but we will demonstrate the idea of the proof on a simple special case, which also makes the statement of the theorem more tangible. In the case of binary classification, $\mathbf{Y}:=\{-1,1\}$, the most natural choice of $B$ is $B(\{y\}\mid(x,y)):=0$, so that the Markov kernel sends every example $(x,y)$ to the other label $-y$. We can then rewrite (7) as
$G(z_{1},\dots,z_{n},z_{n+1}):=\mathrm{e}^{-1}\frac{E(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})},$ (8)
which does not involve any averaging. We can interpret (8) as the conformal e-predictor $E^{\mathrm{X}}$ being almost as efficient as the original train-invariant randomness e-predictor $E$, where efficiency is measured by the degree to which we reject the false label $-y_{n+1}$. For example, for a small positive constant $\epsilon$, $G\geq 1/\epsilon$ with probability at most $\epsilon$, and so
$E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})>\mathrm{e}^{-1}\epsilon E(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})$ (9)
with probability at least $1-\epsilon$.
To get an idea of the proof, suppose the true data sequence $z_{1},\dots,z_{n+1}$, generated in the IID fashion, is such that the ratio in (8) is large. By definition, the ratio $E^{\mathrm{i}}=E/E^{\mathrm{X}}$ is an invariant randomness e-variable, and so only depends on the bag of its input examples and measures the degree to which that bag does not look IID. Flipping the label of a randomly chosen $z_{i}$, $i\in\{1,\dots,n+1\}$, in the training bag $\lbag z_{1},\dots,z_{n+1}\rbag$ leads to a data bag that is still compatible with the assumption of randomness, and if flipping the label $y_{n+1}$ in the training bag leads to a large value of $E/E^{\mathrm{X}}$, this means that the example $z_{n+1}$ was unusual in the training bag, which makes the original data sequence $z_{1},\dots,z_{n+1}$, in which the unusual example is also the last one, not compatible with the assumption of randomness.
The following result is a simple statement of optimality for Theorem 5.
Theorem 6.
The constant $\mathrm{e}^{-1}$ in Theorem 5 cannot be replaced by a larger one, even if we assume $E\in\mathcal{E}^{\mathrm{tR}}$.
In principle, Theorem 6 follows from a later result (namely, Theorem 8 below), but in Sect. A.2 we will give a simple independent proof. The origin of the factor $\mathrm{e}^{-1}$ is the difference between the assumptions of randomness and exchangeability: while flipping the label of a randomly chosen example $z_{i}$, $i\in\{1,\dots,n+1\}$, keeps the exchangeability of the original IID data sequence, the new sequence ceases to be generated in the IID fashion. Therefore, we approximate flipping the label of one randomly chosen example by flipping the label of each example $z_{i}$, $i\in\{1,\dots,n+1\}$, with a small probability; $\mathrm{e}^{-1}$ is the largest probability (which is attainable) that exactly one label will be flipped. Full details are given in Appendix A.
3.2 From train-invariant randomness p-prediction to conformal p-prediction
It is known that, for any $\delta\in(0,1)$, the function $p\mapsto\delta p^{\delta-1}$ transforms p-values to e-values and that the function $e\mapsto e^{-1}$ transforms e-values to p-values. See, e.g., [18, Propositions 2.1 and 2,2]. More generally, any function $f:[0,1]\to[0,\infty]$ integrating to 1 transforms p-values to e-values (is a p-to-e calibrator), while $e\mapsto\min(e^{-1},1)$ is the optimal way of transforming e-values to p-values (it is an optimal e-to-p calibrator). This allows us to adapt Theorem 5 to p-predictors.
Corollary 7.
Let $B:\mathbf{Z}\hookrightarrow\mathbf{Y}$ be a Markov kernel and let $\delta\in(0,1)$. For each randomness p-predictor $P$ there exists an exchangeability p-predictor $P^{\prime}$ such that
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\delta}{\mathrm{e}}\int\frac{P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)}\;B(\mathrm{d}y\mid z_{n+1})$ (10)
is a randomness e-variable.
The proof is obvious (calibrate $P$ to get $E\in\mathcal{E}^{\mathrm{X}}$ and then calibrate $E^{\mathrm{X}}$ to get $P^{\prime}\in\mathcal{P}^{\mathrm{X}}$), but it is still spelled out in Sect. A.3. As before, we concentrate on the case where $P$ is train-invariant, and in this case $P^{\prime}$ can be chosen as conformal p-predictor.
The interpretation of (10) is much simpler in the binary case $\mathbf{Y}=\{-1,1\}$ with the same Markov kernel as before. In this case (10) becomes
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\delta}{\mathrm{e}}\frac{P^{\prime}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})}.$
Therefore,
$P^{\prime}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})<\frac{\mathrm{e}}{\delta\epsilon}P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},-y_{n+1})$
with probability at least $1-\epsilon$.
In conclusion of this section, let us discuss (10) in general. The interpretation of (10) is that, under the randomness of the true data sequence, $P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ is typically small (perhaps not to the same degree) when $P(z_{1},\dots,z_{n},x_{n+1},y)$ is small; i.e., we do not lose much in efficiency when converting train/invariant randomness p-predictors to conformal p-predictors. To see this, fix small $\epsilon_{1},\epsilon_{2}\in(0,1)$. Then we will have $G(z_{1},\dots,z_{n},z_{n+1})<1/\epsilon_{1}$ for the true data sequence $z_{1},\dots,z_{n},z_{n+1}$ unless a rare event (of probability at most $\epsilon_{1}$) happens. For the vast majority of the potential labels $y\in\mathbf{Y}$ we will then have
$\frac{\delta}{\mathrm{e}}\frac{P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}{P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y)}<\frac{1}{\epsilon_{1}\epsilon_{2}},$ (11)
$P^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)<\frac{\mathrm{e}}{\delta\epsilon_{1}\epsilon_{2}}P^{1-\delta}(z_{1},\dots,z_{n},x_{n+1},y),$ (12)
4 Applications to multi-class classification and optimality results
In this section we discuss the case of classification, $\left|\mathbf{Y}\right|<\infty$, but now we are interested in the non-binary case $\left|\mathbf{Y}\right|>2$ (the $\sigma$-algebra on $\mathbf{Y}$ is discrete, as usual). Let us only discuss reduction of randomness e-predictors to conformal e-predictors. Reduction of randomness p-predictors to conformal p-predictors is completely analogous; it just uses (10) instead of (7).
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\mathrm{e}^{-1}}{\left|\mathbf{Y}\right|-1}\sum_{y\in\mathbf{Y}\setminus\{y_{n+1}\}}\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}.$ (13)
As before, we assume that $E$ is train-invariant. The interpretation of (13) is that the conformal e-predictor $E^{\mathrm{X}}$ is almost as efficient as the original train-invariant randomness e-predictor $E$ on average; as before, efficiency is measured by the degree to which we reject the false labels $y\neq y_{n+1}$. Roughly, on average, we lose at most a factor of $\mathrm{e}$ in the e-values of false labels when we replace $E$ by $E^{\mathrm{X}}$.
Of course, we can avoid “on average” by making (13) cruder and replacing it by the existence of $G\in\mathcal{E}^{\mathrm{R}}$ satisfying
$\forall(z_{1},\dots,z_{n})\in\mathbf{Z}^{n}\;\forall x_{n+1}\in\mathbf{X}\;\forall y\in\mathbf{Y}\setminus\{y_{n+1}\}:\\ G(z_{1},\dots,z_{n},z_{n+1})\geq\frac{\mathrm{e}^{-1}}{\left|\mathbf{Y}\right|-1}\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)},$ (14)
where $z_{n+1}:=(x_{n+1},y_{n+1})$. For a small positive constant $\epsilon$, we can then claim that, with probability at least $1-\epsilon$ over the true data sequence,
$\forall y\in\mathbf{Y}\setminus\{y_{n+1}\}:E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)>\frac{\mathrm{e}^{-1}\epsilon}{\left|\mathbf{Y}\right|-1}E(z_{1},\dots,z_{n},x_{n+1},y).$ (15)
$G(z_{1},\dots,z_{n},z_{n+1}):=\frac{\mathrm{e}^{-1}}{\left|\mathbf{Y}\right|}\sum_{y\in\mathbf{Y}}\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}.$ (16)
Let us check that the denominators, $\left|\mathbf{Y}\right|-1$ or $\left|\mathbf{Y}\right|$, in (13), (14), and (16) are asymptotically optimal.
Theorem 8.
For each constant $c>1$ the following statement holds true for a sufficiently large $\left|\mathbf{Y}\right|$ and a sufficiently large $n$. There exists a train-invariant randomness e-predictor $E$ such that for each randomness e-variable $G$ there exist $z_{1},\dots,z_{n}$, $z_{n+1}=(x_{n+1},y_{n+1})$, and $y\neq y_{n+1}$ such that
$G(z_{1},\dots,z_{n},z_{n+1})<\frac{c\mathrm{e}^{-1}}{\left|\mathbf{Y}\right|}\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\mathrm{X}}(z_{1},\dots,z_{n},x_{n+1},y)}.$
When we say “for a sufficiently large $\left|\mathbf{Y}\right|$” in Theorem 8, the lower bound on $\left|\mathbf{Y}\right|$ is allowed to depend on $c$, and when we say “and a sufficiently large $n$”, the lower bound on $n$ is allowed to depend on $c$ and $\left|\mathbf{Y}\right|$.
A complete proof of Theorem 8 is given in Sect. A.4, but the informal idea of the proof is that we can ignore the objects and make a false test label $y\neq y_{n+1}$ encode the bag $\lbag y_{1},\dots,y_{n}\rbag$ of the training labels. We have $E^{\mathrm{i}}:=E/E^{\mathrm{X}}\in\mathcal{E}^{\mathrm{iR}}$. If we also make $y$ easily distinguishable from the training labels $y_{1},\dots,y_{n}$, the value $E^{\mathrm{i}}(z_{1},\dots,z_{n},x_{n+1},y)$ (even if depending only on $\lbag z_{1},\dots,z_{n},x_{n+1},y\rbag$) can be made large, even for a true data sequence generated in the IID fashion.
Formally, Theorem 8 is an inverse to (14) and (16), but it has several weaknesses in this role:
  • It shows that $E^{\mathrm{X}}$ is not competitive with $E$ in some cases, but can there be another conformal e-predictor $E^{\prime}\in\mathcal{E}^{\mathrm{tX}}$ that is better than $E^{\mathrm{X}}$ in this respect?
  • Even if the ratio $E(z_{1},\dots,z_{n},x_{n+1},y)/E^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ for $E^{\prime}\in\mathcal{E}^{\mathrm{tX}}$ is very large, it is not so interesting if already $E^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ is very large.
  • Even if the ratio $E(z_{1},\dots,z_{n},x_{n+1},y)/E^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ for $E^{\prime}\in\mathcal{E}^{\mathrm{tX}}$ is very large, it is not so interesting if $G(z_{1},\dots,z_{n},z_{n+1})$ is very large for the true data sequence.
The following result overcomes these weaknesses.
Theorem 9.
For each constant $c\in(0,1)$ and each label space $\mathbf{Y}$, $1<\left|\mathbf{Y}\right|<\infty$, the following statement holds true for sufficiently large $n$. There exists a train-invariant randomness e-predictor $E$ such that for each conformal e-predictor $E^{\prime}$ and each randomness e-variable $G$ there exist $z_{1},\dots,z_{n}$, $z_{n+1}=(x_{n+1},y_{n+1})$, and $y\neq y_{n+1}$ such that
$G(z_{1},\dots,z_{n},z_{n+1})$ $\leq 2,$ (17)
$E^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)$ $\leq 2.01,$ (18)
$E(z_{1},\dots,z_{n},x_{n+1},y)$ $\geq c\left|\mathbf{Y}\right|.$ (19)
To compare Theorem 9 with Theorem 8, notice that (17)–(19) imply
$G(z_{1},\dots,z_{n},z_{n+1})<\frac{4.02c^{-1}}{\left|\mathbf{Y}\right|}\frac{E(z_{1},\dots,z_{n},x_{n+1},y)}{E^{\prime}(z_{1},\dots,z_{n},x_{n+1},y)}.$
Therefore, Theorem 9 implies Theorem 8, but only if we ignore the constant factor $114.02\mathrm{e}<$ 11.
Proof sketch of Theorem 9.
Let us ignore the objects and set, without loss of generality (assuming our prediction problem is classification), $\mathbf{Y}:=\{0,\dots,\left|\mathbf{Y}\right|-1\}$. Generate labels $Y_{1},\dots,Y_{n+1}$ randomly (independently from the uniform distribution on $\mathbf{Y}$), and set $Y\equiv-Y_{1}-\dots-Y_{n}\ (\mathrm{mod}\,\left|\mathbf{Y}\right|)$; capital letters are used to emphasize that the labels (elements of $\mathbf{Y}$) are random. The idea is to prove that the three events
$G(Y_{1},\dots,Y_{n},Y_{n+1})$ $\leq 2,$ (20)
$E^{\prime}(Y_{1},\dots,Y_{n},Y)$ $\leq 2.01,$ (21)
$E(Y_{1},\dots,Y_{n},Y)$ $\geq c\left|\mathbf{Y}\right|$ (22)
(cf. (17)–(19), respectively) hold with probability at least 0.5, 0.502, and 0.999, respectively. This will imply the statement of the theorem as their intersection will be nonempty.
Since $Y_{1},\dots,Y_{n+1}$ are IID and $Y_{1},\dots,Y_{n},Y$ are exchangeable, the probabilities of (20) and (21) are bounded by Markov’s inequality. And we can define $E$ to ensure that (22) holds with a small probability since $Y_{1},\dots,Y_{n},Y$ are not IID: were they IID, we would typically expect their sum to be divisible by $\left|\mathbf{Y}\right|$ with probability close to $1/\left|\mathbf{Y}\right|$ for a large $n$.
Further details are given in Sect. A.5. ∎
5 Conclusion
This paper gives explicit statements, not involving any unspecified constants, of universality of conformal predictors. Namely, for each train-invariant randomness predictor there is a conformal predictor that is competitive with it. Some constants is these statements have been shown to be optimal.
These are some directions of further research:
  • This paper shows that the attainable improvement on conformal prediction under the assumption of randomness is limited, such as a factor of $\mathrm{e}$ in the e-values for false labels (see, e.g., (8)). Can we develop practically useful predictors exploiting such potential improvements? (For a toy example, see Remark 10 in Sect. A.2.)
  • Can we connect $\mathcal{P}^{\mathrm{tR}}$ and $\mathcal{P}^{\mathrm{tX}}$ directly (in the spirit of Corollary 7), without a detour via e-values?
  • A related remark is that all our optimality results (Theorems 6, 8, and 9) cover only e-prediction. It would be ideal to have similar optimality results for Corollary 7 or its stronger versions.
  • The assumption of randomness is very strong, and there has been extensive work devoted to relaxing this assumption; see, e.g., [13] and [1, Chap. 7]. To what degree do the results of this paper carry over to more general settings?
Acknowledgments
In October 2024 Vladimir V’yugin, one of my teachers in the algorithmic theory of randomness, died at the age of 75. I am deeply grateful to him for his support and encouragement from the time when I first met him, in 1980, until the end of his life. The research reported in this paper is dedicated to his memory.
Appendix A Proofs
This appendix gives detailed proofs of all results stated in the main paper. The Bernoulli model is defined as the statistical model $(B_{\theta}^{n+1}:\theta\in[0,1])$, where $B_{\theta}$ is the Bernoulli measure on $\{0,1\}$, defined by $B_{\theta}(\{1\}):=\theta\in[0,1]$.
A.1 Proof of Theorem 5
@@
Combining all these statements, we get the following randomness e-variable $G^{\prime}$:
$G^{\prime}(z_{1},\dots,z_{n+1})$ $:=G_{2}(z_{1},\dots,z_{n+1})G_{3}(z_{1},\dots,z_{n+1})$
$\geq\mathrm{e}^{-1}G_{1}(z_{1},\dots,z_{n+1})G_{3}(z_{1},\dots,z_{n+1})$
$=\mathrm{e}^{-1}\int E^{\mathrm{i}}(z_{1},\dots,z_{n},x_{n+1},y)B(\mathrm{d}y\mid z_{n+1}).$ (23)
By the definition (6), (7) is equal to the expression in (23), and so $G^{\prime}\in\mathcal{E}^{\mathrm{R}}$ implies that (7) is also a randomness e-variable.
A.2 Proof of Theorem 6
Without loss of generality we assume $\left|\mathbf{X}\right|=1$ (so that the objects become uninformative and we can omit them from our notation) and $\mathbf{Y}=\{-1,1\}$ (with the discrete $\sigma$-algebra). Define a randomness e-variable $E$ by
@@ A.2 Proof of Theorem 6
where $k$ is the number of 1s among $y_{1},\dots,y_{n+1}$. This is indeed a randomness e-variable, since the maximum probability of $k=1$ in the Bernoulli model, $(n+1)p(1-p)^{n}\to\max$, is attained at $p=\frac{1}{n+1}$. The corresponding exchangeability e-variable is
@@
Both $E$ and $E^{\mathrm{X}}$ are train-invariant (and even invariant). Let $B$ just flip the label: $B(\{-y\}\mid y)=1$. Suppose Theorem 5 holds with the $\mathrm{e}^{-1}$ in (7) replaced by $c>\mathrm{e}^{-1}$. Then the randomness e-variable $G$ satisfies
@@
which is impossible for a large enough $n$ (since the probability measure concentrated on $(0,\dots,0)$ is of the form $Q^{n+1}$).